# HRFT: Mining High-Frequency Risk Factor Collections End-to-End via Transformer

### v1

@@ Abstract
In quantitative trading, it is common to find patterns in short-term volatile trends of the market. These patterns are known as High-Frequency (HF) risk factors, serving as key indicators of future stock price volatility. Traditionally, these risk factors were generated by financial models relying heavily on domain-specific knowledge manually added rather than extensive market data. Inspired by symbolic regression (SR), which infers mathematical laws from data, we treat the extraction of formulaic risk factors from high-frequency trading (HFT) market data as an SR task. In this paper, we challenge the manual construction of risk factors and propose an end-to-end methodology, Intraday Risk Factor Transformer (IRFT), to directly predict complete formulaic factors, including constants. We use a hybrid symbolic-numeric vocabulary where symbolic tokens represent operators/stock features and numeric tokens represent constants. We train a Transformer model on the HFT dataset to generate complete formulaic HF risk factors without relying on a predefined skeleton of operators (+, $\times$, ///, $\sqrt{x}$, $\log{x}$, $\cos{x}$). It determines the general shape of the stock volatility law up to a choice of constants, e.g., $f(x)=\tan(ax+b)$ ($x$ is the stock price). We refine the predicted constants (a, b) using the Broyden–Fletcher–Goldfarb–Shanno algorithm (BFGS) to mitigate non-linear issues. Compared to the 10 approaches in SRBench, a living benchmark for SR, IRFT gains a 30% excess investment return on the HS300 and S&P500 datasets, with inference times orders of magnitude faster than theirs in HF risk factor mining tasks.
Introduction
In the past, most research has focused on predicting the expected value of a stock’s future return distribution (Ma, Han, and Wang 2021; Yilmaz and Yildiztepe 2024) by uncovering alphas, also known as return factors (Choi, Jiang, and Zhang 2023). However, high returns often come with high risks. It’s crucial to consider the risk of price volatility to achieve the expected return. (Rezaei, Faaljou, and Mansourfar 2021). A risk factor is a formula that translates historical trading data into a measure of the width of a stock’s future return distribution. As High-Frequency Trading (HFT) markets mature, investors need to respond quickly to short-term market price fluctuations. HFT traders frequently execute orders multiple times during the trading day based on investment managers’ instructions. This paper focuses on capturing short-term volatile trends in HFT markets.
Most risk factor mining relies on experts using financial knowledge to construct models like the Markowitz Mean-Variance Model (Steinbach 2001), CAPM (Merton 1973), and Fama French Models (Taneja 2010; Chiah et al. 2016), then manually selecting the appropriate covariates, also called risk factors. These hand-picked factors often weakly correlate with stock return volatility and lag behind market movements. Statistical models like PCA (Fan, Ke, and Liao 2021; Baek, Cursio, and Cha 2015) and Factor Analysis (Ang, Liu, and Schwarz 2020; Ludvigson and Ng 2007) can obtain latent risk factors but struggle with nonlinear factors. Deep Risk Model (DRM) (Lin et al. 2021) designs risk factors using neural networks to enhance covariance estimation in the Markowitz Mean-Variance Model. However, it focuses on low-frequency risk factors, unsuitable for rapidly changing quantitative trading markets.
@@ Introduction
Inspired by Symbolic Regression (SR), which infers mathematical patterns from large data sets (Landajuela et al. 2022; Becker et al. 2023), we transform raw historical stock data into indicative volatility signals for quantitative trading. HF risk factor mining tasks are essentially SR tasks. We generate formulaic HF risk factors from raw stock trading features to measure short-term market volatility for the first time. These HF risk factors effectively measure the variance of future stock return distributions. We train a Transformer model from scratch on the HFT dataset instead of fine-tuning an open-source Pre-Trained Model (PTM). Note that most of the capabilities of language models lie in semantic processing. However, the task of generating formulas primarily focuses on tackling numbers and symbols, rather than comprehending extensive vocabularies. So to represent formula factors as sequences, we use direct Polish notation. Operators, variables, and integers are single autonomous tokens. Constants are represented as sequences of 3 tokens: sign, mantissa, and exponent. We preprocess open/close/high/low/volume/vwap features using word expressions. Our embedder and transformer learn from word expressions of stock features as input and formulaic factors as output. The sequence-to-sequence Transformer architecture has 16 attention heads and an embedding dimension of 512, totaling 86M parameters.
@@
  • We challenge the manual construction of risk factors and propose a data generator for end-to-end HF risk factor mining for downstream tasks and short-term stock market risk measurements.
  • To do the HF risk factor mining, we employ a hybrid symbolic-numeric vocabulary where symbolic tokens represent operators///stock features and numeric tokens represent constants, then train a Transformer model over the HFT dataset to directly generate full formula factors without relying on skeletons. We refine the prediction constants during the Broyden–Fletcher–Goldfarb–Shanno optimization process (BFGS) as informed guessing, which can alleviate the non-linear issues effectively. Lastly, we utilize generation and inference techniques that enable our model to scale to complex realistic datasets, whereas current works are only suitable for synthetic datasets.
  • Our method outperforms the 10 approaches in SRBench in investment simulation experiments over the HS300 and S&P500 dataset, achieving a 30% higher investment return (refer to Fig. 5) and several orders of magnitude faster inference time in factor mining tasks (refer to Fig. 3).
Figure 1: A formulaic HF risk factor can be represented as a binary tree. The nodes are unary operators like ‘mod’ and ‘abs’, and binary operators like ‘sub’, ‘div’, and ‘mul’. The leaves are input features such as ‘high’, ‘open’, ‘close’, and ‘vwap’, along with constants like 2 and 4.
Related Work
Daily Risk Factors. In the past, risk factors were constructed through traditional factor models, often requiring manual screening of covariates. These covariates, such as beta (Prices 1964), size and value (Fama and French 1992), momentum (Carhart 1997), and residual volatility and liquidity (Sheikh 1996), have several shortcomings. Issues include debates over the best covariate sets (Harvey and Liu 2019), weak correlation with equity return volatility (Anatolyev and Mikusheva 2022; Gospodinov, Kan, and Robotti 2019), and factors lagging behind market changes (Li, Li, and Shi 2017). Additionally, statistical models like principal components (Ait-Sahalia and Xiu 2017; De Nard, Ledoit, and Wolf 2021) and factor analysis (Harman 1976) only yield linear risk factors, which are prone to overfitting (Wan et al. 2023). Furthermore, the deep Risk Model (DRM) (Lin et al. 2021), which combines deep learning with the Markowitz mean-variance model, fails when the loss function is non-trivial.
Symbolic Regression. Symbolic Regression (SR) focuses on identifying mathematical expressions that best describe relationships (Biggio et al. 2021; Kamienny et al. 2022). Mining quantitative factors is essentially an SR task, extracting accurate mathematical patterns from historical trading data. The dominant approach is Genetic Programming (GP) (Lin et al. 2019; Zhang et al. 2020; Cui et al. 2021), but it has two drawbacks: slow execution (Langdon 2011), due to the expansive search space requiring numerous generations to converge, and a tendency towards overfitting (Fitzgerald, Azad, and Ryan 2013), dependent on input fitness. Given the time-sensitive nature of quantitative trading, we aim to generate risk factors that effectively capture short-term market volatility.
Language Models. The pre-trained Finbert(Liu et al. 2021) or other pre-trained Fingpt(Yang, Liu, and Wang 2023; Liu, Wang, and Zha 2023; Zhang, Yang, and Liu 2023) mainly focus on semantic processing. However, the task of generating formulaic factors primarily focuses on tackling numbers and symbols, rather than comprehending extensive vocabularies. They encounter difficulties in complex reasoning tasks, like predicting formulas for large numbers(Mishra et al. 2022). In pretraining, models struggle with rarely or never seen symbols. And they can only deal with sequences. We treat HF risk factors as a language for explaining market volatility trends and train a Transformer. This model offers two benefits: They are fast because of leveraging previous experience with inference being a single forward pass. And they are less over fitting-prone, as they do not require loss optimization based on inputs.
Data generation
@@
Figure 2: Our risk factor-mining framework operates end-to-end. During training, the Embedder concatenates tokens from three sample pairs into a vector, which is then reduced via a two-layer, ReLU-activated Feed-Forward Network (FFN), producing M embedding vectors for the Transformer. In the inference stage, the complete HF risk factor expression is generated directly. We refine the constants in the factor formulas using the BFGS algorithm and incorporate a scaling process during inference to accommodate diverse HF feature sample points.
Generating functions
Inspired by (Lample and Charton 2019), we randomly generate tree structures composed of mathematical operators, variables, and constants, as shown in Fig. 1.
  1. 1.
@@
  3. 3.
    Construct a binary tree using these $b$ binary operator, drawing on the method of (Lample and Charton 2019).
  4. 4.
@@
To independently control the number of unary operators (unrelated to $W$) and binary operators (related to $W$), we cannot directly sample a unary-binary tree as done by (Lample and Charton 2019). We ensure the first $W$ variables are present in the tree, avoiding functions like $x_{2}+x_{4}$ that lack intermediate variables. To enhance expression diversity, we randomly drop out HF variables, allowing our model to set the coefficient of $x_{3}$ to zero. By randomly sampling tree structures and numerical constants, our model rarely encounters identical functions, preventing simple memorization. Detailed parameters for sampling HFT risk factor expressions are in Table 5 (refer to Appendix The detailed parameter set description for more details).
Generating inputs
@@
Tokenization
We preprocess HTF stock open/close/high/low/volume/vwap features using word expressions. Our embedder and transformer learn from these feature word expressions as input and formulaic factor word expressions as output. Following (Charton 2021), we represent numbers in decimal scientific notation with four significant digits, using three tokens: sign, mantissa (0 to $9$), and exponent ($E-100$ to $E100$). Referring to (Lample and Charton 2019), we use prefix expressions (Polish notation) for HF risk factors, encoding operators, variables, and integers, and applying the same method for constants. For example, $f(x)=\tan(9.7341x)$ is encoded as $[tan,mul,+,97341,E-3,x]$. The Decoder’s vocabulary includes symbolic tokens (operators and variables) and integer tokens, while the Encoder’s vocabulary only includes integer tokens.
The Method
@@
Model
In this section, we introduce an Embedder to reduce input dimensions, making IRFT more suitable for high-dimensional transaction data. The Transformer model is sensitive to input samples when predicting factor expressions due to the complementarity of the attention heads. Some heads focus on extremes like $-1,0,1$ of exponential functions, while others concentrate on the periodicity of trigonometric functions.
Embedder
@@
We utilize a sequence-to-sequence Transformer model, as proposed by (Vaswani et al. 2017), with 16 attention head and an embedding dimension of 512, totaling 84 million parameters. Following (Charton 2021), we observe that the solution to the factor mining task is an asymmetric model structure with a deeper decoder. In more details, we use 4 layers in the encoder and 16 layers in the decoder. The encoder effectively captures the distinctive features of HFT stock data, such as periodicity and continuity, by mixing short-ranged attention heads focusing on local features with long-ranged attention heads capturing the global form of the HF risk factor. Unlike previous risk factor mining models (Harman 1976; Lin et al. 2021), our model accounts for the periodicity of stock sample points, capturing short-term volatility and cycling in the HFT stock market.
Training.
@@
We employ the Adam optimizer to minimize the cross-entropy loss. Actually, our cross entropy loss calculates the probability distribution of target tokens and risk factor tokens, in order to find the factor token with the maximum probability of the target token. The learning rate is initially set to $2\times 10^{-7}$ and is linearly increased to $2\times 10^{-3}$ over the first 10,000 steps. Following the approach proposed by (Vaswani et al. 2017), we then decay the learning rate as the inverse square root of the step numbers. To assess the model’s performance, We take 10% samples from the same generator as a validation set and trained IRFT model until the accuracy on the validation set was satisfied. We esitimate this process to span approximately 50 epochs, with each epoch consisting of 3 million sample points. Moreover, to minimize the padding waste, we put together sample point with similar lengths, which can ensure ensure more than 3M samples in each batch.
Inference Tricks
@@
We describe four tricks to enhance IRFT’s inference performance. Previous SR models, like (Biggio et al. 2021; Valipour et al. 2021), use a standard skeleton approach, predicting equation skeletons first and then fitting constants. These methods are less accurate and handle limited input dimensions ($W<3$). Our end-to-end (E2E) method predicts functions and constants simultaneously, allowing our model to process more input features ($W=7$) and handle large real-world datasets (about 26M samples).
Refinement
@@
As suggested by (Kamienny et al. 2022), we significantly enhance effectiveness by adding a refinement step: fine-tuning constants using BFGS with our model predictions as initial values. This improves the skeleton method in two ways: providing stronger supervision signals by predicting complete HF risk factor expressions and enhancing constant accuracy with the BFGS algorithm, thereby boosting overall model performance.
Scaling
@@
At the inference stage, we introduce a scaling procedure to ensure accurate predictive factor formulas from samples with varying means and variances. During training, we scale all HF input points to center their distributions at the origin with a variance of 1. For example, let $E$ be the inferred risk factor, $x$ be the input, $\mu=\text{mean}(x)$, and $\sigma=\text{std}(x)$. Then we replace $x$ with $\bar{x}=(x-\mu)/\sigma$, and the model predicts $\hat{E}(\bar{x})=\hat{E}(\sigma x+\mu)$. This allows us to recover an approximate $E$ by unscaling the variables in $\hat{E}$. This approach makes IRFT insensitive to the scale of input points and represents constants outside $D_{(a,b)}$ as multiplications of constants within $D_{(a,b)}$. Unlike our method, DL-based approaches like (Petersen et al. 2019) often fail when samples fall outside the expected value range.
Bagging and decoding

### v5

@@ Abstract.
In quantitative trading, transforming historical stock data into interpretable, formulaic risk factors enhances the identification of market volatility and risk. Despite recent advancements in neural networks for extracting latent risk factors, these models remain limited to feature extraction and lack explicit, formulaic risk factor designs. By viewing symbolic mathematics as a language—where valid mathematical expressions serve as meaningful ”sentences”—we propose framing the task of mining formulaic risk factors as a language modeling problem. In this paper, we introduce an end-to-end methodology, Intraday Risk Factor Transformer (IRFT), to directly generate complete formulaic risk factors, including constants. We use a hybrid symbolic-numeric vocabulary where symbolic tokens represent operators and stock features, and numeric tokens represent constants. We train a Transformer model on high-frequency trading (HFT) datasets to generate risk factors without relying on a predefined skeleton of operators (e.g., $\times,/,\sqrt{x},\log{x},\cos{x}$). It determines the general form of the stock volatility law, including constants; for example, $f(x)=\tan(ax+b)$, where $x$ is the stock price. We refine the predicted constants ($a,b$) using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm to mitigate non-linear issues. Compared to the ten approaches in SRBench, an active benchmark for symbolic regression (SR), IRFT achieves a 30% higher investment return on the HS300 and S&P500 datasets, while achieving inference times that are orders of magnitude faster than existing methods in HF risk factor mining tasks. Our code and dataset are publicly accessible via the following GitHub repository: https://github.com/wencyxu/IRF-LLM-accepted-at-WWW25-.
Computational Finance, Stock Volatility Forecasting, Transformer; Factor Analysis
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3715235isbn: 979-8-4007-1331-6/25/04ccs: Computing methodologies Transformerccs: Computing methodologies Search methodologiesccs: Applied computing Economics
1. Introduction
Most risk factor mining traditionally relies on factor models and requires the manual screening of covariates, also known as risk factors, such as beta (Prices, 1964), size and value (Fama and French, 1992), momentum (Carhart, 1997), and residual volatility and liquidity (Sheikh, 1996). However, these hand-picked factors often show weak correlations with stock return volatility and fail to keep pace with market dynamics. Statistical models such as Principal Component Analysis (PCA) (Fan et al., 2021) and factor analysis (Ang et al., 2020) can identify latent risk factors but are limited in capturing nonlinear risks. Deep Risk Model (DRM) (Lin et al., 2021) address this limitation by utilizing neural networks to learn and extract latent risk factors directly from stock data, replacing the traditional manually designed covariance matrices. However, when it comes to designing explicit signals—formulaic risk factors representing market trends—neural networks are relatively limited, being primarily restricted to feature extraction.
Symbolic mathematics can be likened to a language, where mathematically valid expressions function as coherent ”sentences.” Inspired by this analogy, we frame the task of extracting formulaic risk factors as a language modeling problem. Thus, employing deep language models to process symbolic mathematics emerges as a natural approach. Inspired by symbolic regression (SR)—which focuses on discovering general mathematical relationships from large datasets (Becker et al., 2023; Xu et al., 2025a)—we aim to uncover patterns underlying the volatility in historical stock data.
@@ 1. Introduction
We generate formulaic HF risk factors from raw stock trading features to measure short-term market volatility for the first time. These HF risk factors effectively measure the variance of future stock return distributions. We train a Transformer model from scratch on the HFT dataset instead of fine-tuning an open-source Pre-Trained Model (PTM). Note that most of the capabilities of language models lie in semantic processing. However, the task of generating formulas primarily focuses on tackling numbers and symbols, rather than comprehending extensive vocabularies. So to represent formula factors as sequences, we use direct Polish notation. Operators, variables, and integers are single autonomous tokens. Constants are represented as sequences of 3 tokens: sign, mantissa, and exponent. We preprocess open/close/high/low/volume/vwap features using word expressions. Our embedder and transformer learn from word expressions of stock features as input and formulaic factors as output. The sequence-to-sequence Transformer architecture has 16 attention heads and an embedding dimension of 512, totaling 86M parameters.
@@
  • We propose a data generator for end-to-end HF risk factor mining for downstream tasks and short-term stock market risk measurements.
  • To do the HF risk factor mining, we employ a hybrid symbolic-numeric vocabulary where symbolic tokens represent operat
    ors//stock features and numeric tokens represent constants, then train a Transformer model over the HFT dataset to directly generate full formula factors without relying on skeletons. We refine the prediction constants during the Broyden–Fletcher–Goldfarb–Shanno optimization process (BFGS) as informed guessing, which can alleviate the non-linear issues effectively. Lastly, we utilize generation and inference techniques that enable our model to scale to complex realistic datasets, whereas current works are only suitable for synthetic datasets.
  • Our method outperforms the 10 approaches in SRBench in investment simulation experiments over the HS300 and S&P500 dataset, achieving a 30% higher investment return (refer to Fig. 5(b)) and several orders of magnitude faster inference time in factor mining tasks (refer to Fig. 3).
Figure 1. A formulaic HF risk factor can be represented as a binary tree. The nodes are unary operators like ‘mod’ and ‘abs’, and binary operators like ‘sub’, ‘div’, and ‘mul’. The leaves are input features such as ‘high’, ‘open’, ‘close’, and ‘vwap’, along with constants like 2 and 4.
Table 1. Detailed parameter settings of IRFT.
Parameter Description Value
W=5(S&P500) W=5(SSE50)
W_max Max input dimension 10
b_max Max binary operators 5+W 5+W
O_b Binary operators add,sub,mul,div
u_max Max unary operators 5 5
O_u Unary operators inv,abs,sqr,sqrt,sin,cos,tan,atan,log,exp
M_min Min number of points 10D
M_max Max number of points 200
c_max Max number of candidate function clusters 10
2. Data generation
@@
2.1. Generating functions
Inspired by (Lample and Charton, 2019), we randomly generate tree structures composed of mathematical operators, variables, and constants, as shown in Fig. 1.
  1. (1)
@@ 2.1. Generating functions
  3. (3)
    Construct a binary tree using these $b$ binary operator, drawing on the method of (Lample and Charton, 2019).
  4. (4)
@@
To independently control the number of unary operators (unrelated to $W$) and binary operators (related to $W$), we cannot directly sample a unary-binary tree as done by (Lample and Charton, 2019). We ensure the first $W$ variables are present in the tree, avoiding functions like $x_{2}+x_{4}$ that lack intermediate variables. To enhance expression diversity, we randomly drop out HF variables, allowing our model to set the coefficient of $x_{3}$ to zero. By randomly sampling tree structures and numerical constants, our model rarely encounters identical functions, preventing simple memorization. Detailed parameters for sampling HFT risk factor expressions are in Table 1.
2.2. Generating inputs
@@
2.3. Tokenization
We preprocess HTF stock open/close/high/low/volume/vwap features using word expressions. Our embedder and transformer learn from these feature word expressions as input and formulaic factor word expressions as output. Following (Charton, 2021), we represent numbers in decimal scientific notation with four significant digits, using three tokens: sign, mantissa (0 to 9999), and exponent (E100E-100 to $E100$). Referring to (Lample and Charton, 2019), we use prefix expressions (Polish notation) for HF risk factors, encoding operators, variables, and integers, and applying the same method for constants. For example, $f(x)=\tan(9.7341x)$ is encoded as [tan,mul,+,97341,E3,x][tan,mul,+,97341,E-3,x]. The Decoder’s vocabulary includes symbolic tokens (operators and variables) and integer tokens, while the Encoder’s vocabulary only includes integer tokens.
Figure 2. Our risk factor-mining framework operates end-to-end. During training, the Embedder concatenates tokens from three sample pairs into a vector, which is then reduced via a two-layer, ReLU-activated Feed-Forward Network (FFN), producing M embedding vectors for the Transformer. In the inference stage, the complete HF risk factor expression is generated directly. We refine the constants in the factor formulas using the BFGS algorithm and incorporate a scaling process during inference to accommodate diverse HF feature sample points.
3. The Method
@@
3.1. Model
In this section, we introduce an Embedder to reduce input dimensions, making IRFT more suitable for high-dimensional transaction data. The Transformer model is sensitive to input samples when predicting factor expressions due to the complementarity of the attention heads. Some heads focus on extremes like 1,0,1-1,0,1 of exponential functions, while others concentrate on the periodicity of trigonometric functions.
3.1.1. Embedder
@@ 3.1.2. Transformer
We utilize a sequence-to-sequence Transformer model, as proposed by (Vaswani et al., 2017), with 16 attention head and an embedding dimension of 512, totaling 84 million parameters. Following (Charton, 2021), we observe that the solution to the factor mining task is an asymmetric model structure with a deeper decoder. In more details, we use 4 layers in the encoder and 16 layers in the decoder. The encoder effectively captures the distinctive features of HFT stock data, such as periodicity and continuity, by mixing short-ranged attention heads focusing on local features with long-ranged attention heads capturing the global form of the HF risk factor. Unlike previous risk factor mining models (Harman, 1976; Lin et al., 2021), our model accounts for the periodicity of stock sample points, capturing short-term volatility and cycling in the HFT stock market.
3.2. Training.
@@
We employ the Adam optimizer to minimize the cross-entropy loss. Actually, our cross entropy loss calculates the probability distribution of target tokens and risk factor tokens, in order to find the factor token with the maximum probability of the target token. The learning rate is initially set to $2\times 10^{-7}$ and is linearly increased to $2\times 10^{-3}$ over the first 10,000 steps. Following the approach proposed by (Vaswani et al., 2017; Xiang et al., 2025), we then decay the learning rate as the inverse square root of the step numbers. To assess the model’s performance, We take 10% samples from the same generator as a validation set and trained IRFT model until the accuracy on the validation set was satisfied. We esitimate this process to span approximately 50 epochs, with each epoch consisting of 3 million sample points. Moreover, to minimize the padding waste, we put together sample point with similar lengths, which can ensure ensure more than 3M samples in each batch.
3.3. Inference Tricks
@@
We describe four tricks to enhance IRFT’s inference performance. Previous SR models, like (Biggio et al., 2021; Valipour et al., 2021), use a standard skeleton approach, predicting equation skeletons first and then fitting constants. These methods are less accurate and handle limited input dimensions ($W<3$). Our end-to-end (E2E) method predicts functions and constants simultaneously, allowing our model to process more input features ($W=7$) and handle large real-world datasets (about 26M samples).
3.3.1. Refinement
@@
As suggested by (Kamienny et al., 2022), we significantly enhance effectiveness by adding a refinement step: fine-tuning constants using BFGS with our model predictions as initial values. This improves the skeleton method in two ways: providing stronger supervision signals by predicting complete HF risk factor expressions and enhancing constant accuracy with the BFGS algorithm, thereby boosting overall model performance.
3.3.2. Scaling
@@
At the inference stage, we introduce a scaling procedure to ensure accurate predictive factor formulas from samples with varying means and variances. During training, we scale all HF input points to center their distributions at the origin with a variance of 1. For example, let $E$ be the inferred risk factor, $x$ be the input, $\mu=\text{mean}(x)$, and $\sigma=\text{std}(x)$. Then we replace $x$ with $\bar{x}=(x-\mu)/\sigma$, and the model predicts $\hat{E}(\bar{x})=\hat{E}(\sigma x+\mu)$. This allows us to recover an approximate $E$ by unscaling the variables in $\hat{E}$. This approach makes IRFT insensitive to the scale of input points and represents constants outside $D_{(a,b)}$ as multiplications of constants within $D_{(a,b)}$. Unlike our method, DL-based approaches like (Petersen et al., 2019) often fail when samples fall outside the expected value range.
3.3.3. Bagging and decoding